OpenStudy (anonymous):

Carrie deposited $5,288.73 into a savings account with an interest rate of 2.3% compounded quarterly. About how long will it take for the account to be worth $8,000? 18 years, 10 months 18 years, 1 month 16 years, 5 month 18 years, 2 months

OpenStudy (mary.rojas):

A formula for calculating annual compound interest is A = P \left(1 + \frac{r}{n}\right)^{nt} where A = value after t periods P = principal amount (initial investment) r = annual nominal interest rate (not reflecting the compounding) n = number of times the interest is compounded per year t = number of years the money is borrowed for

OpenStudy (anonymous):

\[8000 = 5288.73( 1+\frac{ i }{ 4 })^{4x}\] solve for x

OpenStudy (mary.rojas):

plug in what you know and solve the missing letter

OpenStudy (mary.rojas):

@mebs thats how :)

OpenStudy (anonymous):

AHHHH yea =)) @mary.rojas

OpenStudy (mary.rojas):

@mebs isnt there something special you have to do to solve for x in this case

OpenStudy (mary.rojas):

like ln or log or something

OpenStudy (anonymous):

So D?

OpenStudy (anonymous):

Ohh yea we could do that good idea but \[\sqrt[4]{ }\] is a better idea...

OpenStudy (anonymous):

Actually your right @mary.rojas you have to do ln1.51265 =4xln(whatever)

OpenStudy (anonymous):

Wait no, it's 16 years and 2 months. That's what I got with the formula you said

OpenStudy (anonymous):

leme check

OpenStudy (mary.rojas):

yeahh i forgot how to do that. but you can plug each choice in for x until you get the desired amount.

OpenStudy (anonymous):

Its 18.04

OpenStudy (anonymous):

18.04? That isn't a choice

OpenStudy (anonymous):

its in years

OpenStudy (anonymous):

18 years and 4 months??? That still isn't a choice

OpenStudy (anonymous):

Not 4 months

OpenStudy (anonymous):

its 18 years and 1 month....

OpenStudy (anonymous):

You round it... 0.0457333 * 12 = 0.5476 = ~~1....

OpenStudy (mary.rojas):

1 month divided by 12 months equals 0.08333 so 18 years and 1 month is equal to 18.08333333

OpenStudy (anonymous):

Just pick the closest one...

OpenStudy (mathstudent55):

\(A = P \left(1 + \frac{r}{n}\right)^{nt}\) \(8000 = 5288.73 \left(1 + \dfrac{0.023}{4}\right)^{4t}\) \(1.51265 = \left(1.00575 \right)^{4t}\) \( \log {1.51265} = \log{\left(1.00575 \right)^{4t} }\) \(0.179739 = 4t \log{1.00575 }\) \(72.18297 = 4t\) t = 18.05 ~years \( 0.05 ~years \times \dfrac{12 ~months}{1 ~year} = 0.55 ~months \approx 1 ~month\) Answer: 18 years and 1 month

OpenStudy (mary.rojas):

yayyy! @mathstudent55

OpenStudy (anonymous):

That's what i am talking about!! @mathstudent55 @mary.rojas

OpenStudy (anonymous):

Thank you all! Wish I could give 3 medals, but I can't haha

OpenStudy (mary.rojas):

:)

OpenStudy (anonymous):

your welcome =))

OpenStudy (anonymous):

Do any of you know how to find monthly payments? I'm making a separate question for it, if you know how please help! x.x

OpenStudy (mathstudent55):

@mebs did it correclty and before me. He deserves a lot of credit and the medal I gave him.

OpenStudy (anonymous):

the monthly payment formula.. \[P = R \frac{[( 1-(1+\frac{ i }{ n })^{-nt} ]}{ \frac{ i }{ n } }\]

OpenStudy (anonymous):

Thanks haha @mathstudent55

OpenStudy (anonymous):

You can also use.. \[M= \frac{ P \times i }{ (c)[(1-(1+\frac{ i }{ c) })]^{-nc} }\]

OpenStudy (anonymous):

I'm not sure what I'm doing wrong, I've plugged everything in but i'm getting a large number

OpenStudy (anonymous):

well how large is it? and what are you investing is it compounded yearly and what were you expecting... Best thing to do is to repost this as a new question...

OpenStudy (anonymous):

Rita owes $9,739 on a credit card with a 21.5% interest rate compounded monthly. What monthly payment should she make to pay off this debt in six years, assuming she does not charge any more purchases with the card?

OpenStudy (anonymous):

Is $20.15 an option. ...

OpenStudy (anonymous):

No

OpenStudy (anonymous):

$909.17 $135.26 $241.82 $2,093.89

OpenStudy (anonymous):

its 241082

OpenStudy (anonymous):

$241.82 yea my bad not 241082

OpenStudy (anonymous):

\[\frac{ 9739 \times 0.215 }{[ (12)(1-(1+0.017916666)^{-72}] }\] try that on ur calculator

OpenStudy (anonymous):

8.6588 ?? Idk

OpenStudy (anonymous):

no try again.. you should get 241.82

OpenStudy (mathstudent55):

\( M = \dfrac{rP}{1 - (1 + r)^{-n} } \) where M = monthly payment r = annual rate divided by 12 and expressed as a decimal n = number of months \( M = \dfrac{0.01791667 \times $9739}{1 - (1 + 0.01791667)^{-72} } \) \( M = \dfrac{$174.4904}{1 - 0.2784324 } \) \( M = \dfrac{$174.4904}{0.7215676} \) \(M = $241.82 \) Answer is C.